Order: Cefaclor (Ceclor) oral suspension 50mg/kg/day in 4 divided doses. Child weighs 22 lb. Available: Cefaclor 250 mg/5mL When instructing the parents to administer Cefaclor, the nurse will tell them to administer _____ tsp for each dose. ________ tsp

pls show all work, tu
Answer #1

ANSWER.

* Calculation of amount(tsp)of drug required per dose.

        Formula = Desired dose/Dose in hand * Quantity.

     Here

         Ordered dose = 50 mg/kg/day in 4 divided doses.

         Weight of the patient = 22 lb.

    Convert lb to kg for calculation.

                1 lb = 0.45 kg.

   So

       Weight of the patient = 22 * 0.45

                                         = 9.9 kg

                                         = 10 kg.

       Desired dose(D) = 50 * 10

                                   = 500 mg.

       Dose in hand(H) = 250 mg.

           Quantity(Q) = 5 ml.

As per formula

       Amount of drug required = 500/250 * 5

                                               = 2 * 5

                                              = 10 ml / day four divided doses.

      Amount(ml) of drug required per dose = 10/4

                                                                    = 2.5 ml.

                           1 tsp(teaspoon) = 5ml.

      So

           Amount(teaspoon) of drug required per dose = 2.5/5

                                                                              = 0.5 teaspoon.